OpenStudy (bossimbacon):

Write your answer in scientific notation The state of Wyoming is roughly the shape of a rectangle that is 3.6 x 10^2 mi long and 2.8 x 10^2 mi wide. What is the approximate area of Wyoming? Hint: The area of a rectangle is the product of its length and width. @thatgirlcaleigh

OpenStudy (anonymous):

100,800

OpenStudy (bossimbacon):

@thatgirlcaleigh wow your really smart and pretty<3 thank you

OpenStudy (anonymous):

It's not too difficult. I'm in Honor's Pre-Calc so I've had a lot of experience with Math. :)

OpenStudy (anonymous):

caleigh!!!

OpenStudy (campbell_st):

well the answer needs to be in scientific notation

OpenStudy (anonymous):

stop giving full answers!!

OpenStudy (anonymous):

yeah and their wrong 2

OpenStudy (campbell_st):

so perhaps reading the question is a skill you need to develop to be an honors student

OpenStudy (anonymous):

yes indeed

OpenStudy (anonymous):

LOL. Okay. Then put it in scientific notation. It's not hard.

OpenStudy (anonymous):

Move the decimal place until you have a number between 1 and 10. Then you add a power of ten that tells how many places you moved the decimal.

OpenStudy (anonymous):

There, not telling the full answer. ;)

OpenStudy (bossimbacon):

1.00x10 to the 5th.?

OpenStudy (bossimbacon):

@thatgirlcaleigh

OpenStudy (anonymous):

1.008000e+1

OpenStudy (bossimbacon):

thanks!

OpenStudy (anonymous):

1.008000 x 10^1

OpenStudy (anonymous):

If that makes sense.

OpenStudy (e.mccormick):

You did your notation wrong. It is not \(10^1\)

OpenStudy (campbell_st):

well if you do it correctly \[3.6 \times 10^2 \times 2.8 \times 10^2 = 1.008 \times 10^5\] you can input it into a calculator as its written... just makesure you display is in Sci mode so that the answer will be given in scientific notation

OpenStudy (anonymous):

That is right. I guess I fluttered that one up. lol

OpenStudy (anonymous):

Been awhile since I've done scientific notation.

OpenStudy (campbell_st):

well the good thing about scientific notation is that it can be done on a calculator, just ensure the display is in scientific mode so that the answer will be given using scientific notation.